OpenStudy (anonymous):

Describe the steps you would use to solve the equation 5(2a - 3)=20 + a

OpenStudy (anonymous):

Use PEMDAS

OpenStudy (anonymous):

5*2a= 10a 5*-3= -15 10a-15= 20+a 10a-a= 9a 9a-15=20 -15 + 20= 5 9a=5 9/9a=a 9/5= 1.8 answer 1.8

OpenStudy (anonymous):

@smileymichelle1996 got the answer

OpenStudy (anonymous):

yes he did

OpenStudy (anonymous):

im a she by the way the one on the left of the picture